### v1

# One-Step Diffusion-Based Image Compression with Semantic Distillation

Abstract
While recent diffusion-based generative image codecs have shown impressive performance, their iterative sampling process introduces unpleasing latency. In this work, we revisit the design of a diffusion-based codec and argue that multi-step sampling is not necessary for generative compression. Based on this insight, we propose OneDC, a One-step Diffusion-based generative image Codec—that integrates a latent compression module with a one-step diffusion generator. Recognizing the critical role of semantic guidance in one-step diffusion, we propose using the hyperprior as a semantic signal, overcoming the limitations of text prompts in representing complex visual content. To further enhance the semantic capability of the hyperprior, we introduce a semantic distillation mechanism that transfers knowledge from a pretrained generative tokenizer to the hyperprior codec. Additionally, we adopt a hybrid pixel- and latent-domain optimization to jointly enhance both reconstruction fidelity and perceptual realism. Extensive experiments demonstrate that OneDC achieves SOTA perceptual quality even with one-step generation, offering over 40% bitrate reduction and 20$\times$ faster decoding compared to prior multi-step diffusion-based codecs. Code will be released later.
footnotetext: Naifu Xue and Zhaoyang Jia are visiting students at Microsoft Research Asia. This paper stems from an open-source project.
1 Introduction
The rapid growth in image data has led to increased storage and transmission costs, heightening the need for efficient, high-quality image compression techniques. While VAE-based learned image compression (LIC) [31, 63] has surpassed traditional codecs (e.g., VVC [6]) in rate-distortion (RD) performance, it often produce blurry details at low bitrates [36]. To mitigate this, some approaches [36, 39] focus on optimizing visual quality by incorporating perceptual metrics and adversarial losses to enable generative image compression. Despite these advances, such generative models tend to introduce artifacts under extreme compression, degrading the realism of reconstructed images.
Recently, diffusion-based generative codecs [7, 28] have been introduced to enhance reconstruction quality by leveraging the powerful content synthesis capabilities of pretrained models. While these methods significantly improve perceptual realism, they may occasionally generate content that deviates from the original input, thereby compromising reconstruction fidelity. In addition, their inherently iterative sampling process leads to substantial computational overhead, making them notably slower than conventional VAE-based codecs.
In standard diffusion-based image generation, the model progressively refines a noisy signal through iterative denoising, beginning with coarse structures and gradually synthesizing high-frequency details [57]. However, when applied to image compression, the task changes: at low bitrates, codecs typically retain coarse structures in the compressed latents. As a result, the decoder is primarily responsible for plausibly reconstructing high-frequency content from the preserved low-frequency information. This observation motivates our central hypothesis: given the compressed latent, multi-step sampling is not necessary for decoding, and a carefully designed one-step alternative could suffice. Although recent advances in one-step diffusion [58, 20, 11] offer promising tools for accelerating inference, their potential remains underexplored in the context of image compression.
Figure 1: Top: multi-step sampling is not essential for image compression; intermediate results are from DiffEIC [28]. Bottom: Visual comparisons including existing open-sourced multi-step diffusion codecs [7, 28, 49] and our proposed one-step codec. Our method achieves the highest visual quality at the lowest bitrate while offering significantly faster decoding.
In this paper, we introduce OneDC (One-step Diffusion-based generative Codec), a novel framework for ultra-low bitrate image coding. OneDC integrates a latent compression module with a one-step diffusion generator: the former encodes the image into compact latent, while the latter synthesizes high-frequency details conditioned on the latent. Since the pixel-domain training is insufficient for guaranteeing perceptual quality at low bitrates [18], we adopt a hybrid-domain training strategy. Specifically, we combine a pixel-domain perceptual loss to promote fidelity with a latent-domain diffusion distillation objective [58] to enhance realism. This design allows OneDC to effectively leverage the pretrained diffusion model while achieving a balanced trade-off among compression ratio, perceptual quality, and decoding efficiency, as illustrated in Fig. 1.
Moreover, we further explore the role of semantic guidance (i.e., the input of the cross-attention layers [44]) within our framework. In one-step diffusion, such guidance is essential to compensate for the absence of multi-step refinement (See ablation in Section 4.3). While existing diffusion models have utilized text prompts for conditioning, we argue that text is suboptimal for image compression due to two limitations: (1) natural language struggles to capture fine-grained or localized visual semantics, and (2) generating high-quality captions typically requires large-scale vision-language models (e.g., the large BLIP2 [27] used in [7]), introducing substantial computational overhead. This raises a central question: Can we design more effective semantic representations than text for guiding one-step diffusion-based codecs—without incurring excessive computation cost?
Recent studies have shown that the hyperprior in VAE-based codecs can capture high-level semantic information [18]. Compared to textual prompts, hyperpriors provide more precise and spatially aligned semantic cues, making them well-suited for representing localized semantics in high-resolution images (see Fig.2). This insight motivates us to further enhance the representational capability of the hyperprior. Inspired by generative tokenizers [12], where codebooks encode rich semantic content, we propose hyperprior semantic distillation—a training strategy that transfers semantic knowledge from a pretrained tokenizer to the hyperprior through a code prediction module [64]. As illustrated in Fig.2, this approach results in reconstructions that are more semantically accurate and visually coherent.
Experiments show that OneDC achieves state-of-the-art (SOTA) performance in generative image compression, delivering over 40% bitrate reduction and 20$\times$ faster decoding compared to existing multi-step diffusion codecs. Ablation studies further validate the effectiveness of our optimization strategy and the semantic distillation method for hyperprior. Our contributions are as follows:
Figure 2: Reconstructions from different semantic guidance. (a) Text prompts (from GPT-4o [41]) struggle to capture complex visual semantics, resulting in severe distortions when using a pretrained text-to-image one-step diffusion model [58]. (b) We finetune the model [58] for hyperprior-to-image generation. Hyperprior guidance yields more faithful reconstructions. (c) Our proposed semantic distillation further improves object-level accuracy, particularly in the highlighted regions.
  • We propose OneDC, a one-step diffusion codec comprising a latent compression module for compact feature encoding and a one-step diffusion generator for fast, high-quality decoding. A hybrid-domain training strategy further enhances both fidelity and perceptual realism.
  • We identify the importance of high-level semantic guidance in one-step diffusion and highlight the limitations of text embeddings. To improve reconstruction quality, we introduce hyperprior features as an alternative and enhance them through semantic distillation.
  • Extensive experiments show that OneDC achieves SOTA compression performance while offering significantly faster decoding than existing diffusion-based codecs, demonstrating the potential of one-step diffusion in generative compression.
2 Related Work
VAE-based Generative Image Compression. VAE-based codecs [31, 63, 50, 4, 38] often produce blurry reconstructions at low bitrates, limiting their perceptual quality. To overcome this, GAN-based approaches such as Agustsson et al. [2] and HiFiC [36] train models to synthesize textures using adversarial discriminators. MS-ILLM [39] improves realism through a local likelihood model, while EGIC [22] introduces semantic-aware feedback via segmentation-conditioned discriminators. Other methods enhance compression by integrating generative tokenizers: Mao et al. [35] utilize VQGAN [12] for compression, Xue et al. [55] unify token generation and entropy coding, and Jia et al. [18] propose GLC that improves token compression via transform coding. More recently, dual-branch codecs [34, 56] have been introduced to support more compact and flexible latent representations. While these methods highlight the strong feature learning capacity of VAE-based codecs, their reconstruction quality at extremely low bitrates remains limited by model capacity and training scale.
Diffusion-based Generative Image Compression. Recent works [24, 7, 28, 49, 54] have explored the use of generative priors from large-scale pretrained diffusion models in image compression, demonstrating improved perceptual realism compared to conventional VAE-based approaches. Eric et al. [24] combinepretrained diffusion with ControlNet [61] to reconstruct images from captions and compressed sketches. PerCo [7] fine-tunes diffusion models using captions and quantized features, with captions generated by a large BLIP2 model [27]. DiffEIC [28] conditions diffusion models on VAE latents, showing strong performance even without text, and their subsequent work [29] improves encoder and introduces relay fine-tuning. Relic et al. [43] treat quantization error as noise and apply denoising diffusion for reconstruction. DiffC [49] demonstrates zero-shot compression with diffusion models using reverse-channel coding. While these approaches improve realism at ultra-low bitrates, they often struggle to maintain high fidelity. Moreover, multi-step sampling in DDIM [45] introduces substantial decoding overhead, limiting their practicality. By contrast, our OneDC achieves high fidelity and realism with significantly faster decoding via one-step sampling. While prior work [28] suggests high-level semantics (e.g., text) may be optional, we find such guidance remains crucial in our one-step diffusion for image decoding task, as detailed in Section 4.3.
One-step Diffusion Models. To reduce the high computational cost of diffusion models, recent efforts have focused on distilling multi-step models into efficient one-step generators. Yin et al. [59] propose Distribution Matching Distillation (DMD), which minimizes an approximate KL divergence between real and generated image distributions modeled by multi-step teacher networks. Their subsequent work, DMD2 [58], further improves generation quality by introducing a diffusion-GAN framework and enhanced training strategies. Based on DMD, Song et al. [46] introduce class-specific student models to improve performance across diverse categories. SwiftBrush series [40, 9] adopt a LoRA adaptation [17] for variational score distillation and further optimize diversity by integrating a clamped CLIP loss. Independently, Kang et al. [20] employ perceptual loss supervision and a multi-scale U-Net discriminator for one-step distillation. These models have also been extended to low-level vision tasks such as super-resolution [11, 53, 51, 21] and image restoration [15], highlighting the potential of one-step diffusion beyond generation. Inspired by these advances, we explore its application in generative image compression, a setting that remains underexplored.
3 Methodology
Figure 3: Overview of the OneDC framework. Q denotes scalar quantization, and FSQ stands for finite scalar quantization. AE and AD refer to the arithmetic encoder and decoder, respectively. $h_{ctx}$ and $h_{sem}$ represent the context and semantic decoders used in the hyperprior branch.
We propose OneDC, a one-step diffusion-based generative image compression framework that achieves high-quality reconstruction with low-latency decoding under ultra-low bitrate constraints. Section 3.1 introduces the coding pipeline of our method, consisting of a latent compression module that encodes the input image into compact latent representation and a one-step diffusion generator that synthesizes reconstructions with realistic details. Given the critical role of semantic guidance in our one-step diffusion and the limitations of textual prompts in compression scenarios, Section 3.2 introduces the use of the hyperprior as an alternative conditioning signal. To further tap the potential of the hyperprior, we propose a semantic distillation strategy that transfers knowledge from a pretrained generative tokenizer to the hyperprior codec, thereby improving semantic accuracy. Finally, Section 3.3 presents our training scheme, which combines pixel-domain and latent-domain objectives to jointly optimize reconstruction fidelity and perceptual realism.
3.1 Framework Overview
Fig. 3 shows the overview of our OneDC framework. On the encoder side, an analysis transform encodes the input $x$ into a compact latent representation $y=g_{a}(x)$, which is then quantized to $\hat{y}$. A hyper encoder further processes the latent into a hyperprior $z=h_{enc}(y)$, quantized as $\hat{z}$. Then, the entropy model takes decoded hyperprior context $c_{h}=h_{ctx}(\hat{z})$ and spatial context $c_{s}$ to predict the distribution of $\hat{y}$, enabling bitrate estimation during training and entropy coding during inference.
Figure 4: Two stage training pipeline of OneDC. The codebook in semantic distillation is initialized from the pretrained tokenizer, and the discriminator in diffusion distillation is abbreviated as Disc.
On the decoder side, the received $\hat{y}$ is passed through a synthesis transform to produce an initial latent $\tilde{y}_{t}=g_{s}(\hat{y})$ for diffusion generation. Given the distinct roles of entropy modeling and semantic processing—where the former estimates low-dimensional distribution parameters and the latter represents high-dimensional visual contents—we introduce an additional semantic decoder for hyperprior to better address their different requirements. It extracts semantic guidance $c=h_{sem}(\hat{z})$ from the quantized hyperprior, which is injected into the cross-attention layers of the one-step diffusion generator. The diffusion generator refines the initial latent in only one step: $\tilde{y}_{0}=\epsilon_{\theta}(\tilde{y}_{t},c)$, which is then decoded by a pretrained VAE decoder to produce the final reconstruction $\hat{x}=D(\tilde{y}_{0})$.
3.2 Semantic Guidance with Hyperprior
3.2.1 From Text to Hyperprior
In the one-step diffusion setting, semantic guidance (the input of the cross-attention layers [44]) plays an even more critical role than in conventional multi-step diffusion. Unlike multi-step methods, which iteratively refine the output and can gradually correct semantic inconsistencies, one-step diffusion models rely entirely on a single forward pass—making accurate semantic conditioning essential. While existing diffusion models typically rely on textual prompts to provide high-level semantic guidance, such prompts are suboptimal in the context of natural image compression. Text struggles to capture localized semantics and often fails to describe fine-grained visual attributes—such as object boundaries and textures—particularly in high-resolution imagery. In addition, generating text descriptions typically requires large models [27], introducing significant computational overhead.
Recent work [18] suggests that the categorical distributed hyperprior with vector quantization (VQ) can capture semantic information. Building on this insight, OneDC replaces textual prompts with categorical hyperprior $\hat{z}$ as semantic guidance for the diffusion model. Specifically, we adopt finite scalar quantization (FSQ) [37], a simpler and more effective quantization scheme than VQ [12], to learn a categorical $\hat{z}$ with semantic information. The hyperprior $z$ is compressed into 7 channels with 4 quantization levels each, yielding an equivalent codebook size of 16,384. At a 64×64$\times$64 × spatial downsampling, this design results in a bitrate of just 0.0034 bpp, incurring very low transmission overhead. Our method offers three key advantages: (1) As shown in Fig. 2, conditioning the one-step diffusion model on categorical hyperprior tokens yields more accurate reconstructions than using text prompts, validating the effectiveness of hyperprior-derived semantic guidance. (2) The 64$\times$ downsampled hyperprior has a large receptive field while preserving spatial locality, effectively balancing global-local semantic expression. (3) Unlike separately trained text encoders [42], our approach enables end-to-end optimization of semantic guidance, allowing the hyperprior to adapt jointly with the diffusion model for improved content alignment.
3.2.2 Semantic Distillation for Hyperprior
While the hyperprior effectively captures semantic cues for diffusion guidance, its representation capability can be further improved. In particular, we propose a semantic distillation mechanism aimed at fully tapping the potential of the hyperprior branch. This distillation guides the hyperprior encoder to better capture visual content and improve the decoding quality of the semantic decoder $h_{sem}$. Specifically, we transfer knowledge from a pretrained high-capacity generative tokenizer [12] to the hyperprior codec via an auxiliary distillation task. This choice is motivated by the structural similarity between the categorical hyperprior codec and the generative tokenizers that are known to produce semantically rich discrete representations [60]. By leveraging this compatibility, we use the tokenizer as a teacher to guide the hyperprior toward more effective semantic encoding.
Following GLC [18], we introduce a transformer-based predictor $P_{aux}$ [64], trained to predict discrete token labels from the hyperprior-derived semantics $c$. The ground truth labels $I_{gt}$ are obtained from a pretrained tokenizer encoder $E_{aux}$. Since the tokenizer is optimized for perceptual reconstruction, its objectives align well with OneDC, making it an effective teacher. Moreover, the smaller information bottleneck in the hyperprior branch naturally filters out redundant information, distilling the most salient semantic features from the teacher. The distillation is supervised via cross-entropy (CE) loss:
$I_{gt}=VQ(E_{aux}(x)),\;\;\;\;logits=P_{aux}(c),\;\;\;\;L_{aux}=CE(I_{gt},logits)$ (1)
Both $P_{aux}$ and $E_{aux}$ are used only during training, introducing no inference overhead. As shown in Fig.2, our semantic distillation enhances content fidelity, and ablation results in Section4.3 further validate its effectiveness in improving final reconstruction quality.
Figure 5: Rate-distortion comparisons. See supplementary for more results.
3.3 Training Strategy
To enhance training efficiency and stability, we adopt a two-stage training strategy, inspired by prior works [36, 39]. As shown in Fig. 4, the stage I mainly trains the compression module while the stage II focuses on fine-tuning the one-step diffusion model for better reconstruction quality. In both stages, the one-step diffusion model is adapted via LoRA layers [17], allowing fast convergence for the compression task while preserving rich generative priors.
Stage I: Pixel-domain Compression Learning This stage aims to: (1) train the codec to compress images into compact latent representations with high-fidelity reconstruction, (2) embed semantic information into the hyperprior via distillation, and (3) adapt the one-step diffusion model to synthesize fine-grained details on the decoder side. The total loss is defined as:
$L_{stageI}=L_{recon}+\lambda R+\alpha L_{aux},\;\;\;\;\text{where}\;\;L_{recon}=L_{1}+L_{perceptual}$ (2)
We use LPIPS [62] as the perceptual loss and $L_{1}$ as the pixel loss. $R$ is the bitrate loss from the quad-tree-based spatial entropy model [25, 26], and $L_{aux}$ is the semantic distillation loss (Eq. 1). The hyperparameter $\lambda$ controls the compression ratio and $\alpha$ is the weight of semantic distillation loss.
Stage II: Hybrid-domain Perceptual Learning. This stage fine-tunes the one-step diffusion model to enhance perceptual realism without degrading fidelity. The latent compression module is fixed to improve training stability [36]. Prior work has shown that pixel-level optimization alone is insufficient for optimal perceptual quality [18]. Thus, we adopt a joint training strategy in both pixel and latent domains. Specifically, we integrate the diffusion-based distillation method from [58], which transfers generation knowledge from a pretrained multi-step diffusion teacher. The training also involves both perceptual supervision from the pixel domain and adversarial alignment in the latent domain, which ensures the fidelity by aligning reconstruction with the original image. The total loss is:
$L_{stageII}=L_{distill}+\beta L_{recon}+\gamma L_{adv},\;\;\;\;\text{where:}$: (3)
$L_{distill}=\mathbb{E}_{t,\tilde{y}^{\prime}_{t}}[\epsilon_{fake}(\tilde{y}^{\prime}_{t},t)-\epsilon_{real}(\tilde{y}^{\prime}_{t},t)],\;\;L_{adv}=\mathbb{E}_{t,\tilde{y}^{\prime}_{t}}[-Disc(\epsilon_{fake}(\tilde{y}^{\prime}_{t},t),t)]$ (4)
Here, $\epsilon_{real}$ and $\epsilon_{fake}$ denote the real and fake score network in the teacher model, and $\tilde{y}^{\prime}_{t}$ is the noised latent at timestep $t$. $Disc$ refers to the discriminator. $\beta$ and $\gamma$ balance the pixel and adversarial losses. Additional implementation details are provided in the supplementary material.
4 Experiment
Figure 6: Visual examples on the CLIC2020 test set. Zoom in for better view.
4.1 Implementation and Evaluation Settings
Model and Training. Our generator adopts the U-Net architecture from the Stable Diffusion 1.5 (SD1.5) [44] and is initialized with weights from a pretrained one-step text-to-image task [58]. In Stage I, we use the pretrained tokenizer from MaskGIT [8] for semantic distillation. In Stage II, we employ the multi-step SD1.5 pretrained model as the teacher for diffusion distillation. To improve high resolution adaptability, we randomly crop patches of size 512 or 1024 during training. Models are optimized using AdamW [33]. Additional settings are provided in the supplementary material.
Evaluation datasets and metrics. We evaluate OneDC on several datasets, including Kodak [13], CLIC2020 test set [47], and MS-COCO 30K [30]. Reconstruction fidelity is assessed using reference perceptual metrics LPIPS [19] and DISTS [10], while generation realism is measured using the no-reference perceptual metric FID [16]. Bitrate saving is measured by BD-Rate [5]. It is worth noting that FID is not reported on the Kodak dataset due to its small size (24 images), which makes reliable FID computation infeasible [7]. More results are provided in the supplementary material.
Compared methods. We focus on comparing our OneDC with generative image codecs at ultra-low bitrate scenarios. The compared methods include the vanilla VAE-based generative codec MS-ILLM [39] and GLC [18], and the recent diffusion-based methods: PerCo [7], DiffEIC [28] and DiffC [49] (SD2.1 based). All methods are evaluated on the full resolution. Given that some diffusion-based codecs are evaluated on the CLIC2020 dataset with cropped $768\times 768$ resolution [7], we also provide the results on this setting in our supplementary material for better comparison.
4.2 Main Results
Quantitative Evaluation. Fig. 9 presents a quantitative comparison between OneDC and existing methods across multiple distortion metrics. Across all datasets—including high-resolution CLIC2020, low-resolution Kodak and MS-COCO 30K—OneDC consistently achieves the best reconstruction fidelity, as measured by LPIPS and DISTS. Compared with the previous SOTA multi-step diffusion-based DiffC, OneDC achieves a 50.08% bitrate saving on Kodak and 56.02% on CLIC2020 in terms of LPIPS quality, demonstrating its superior compression efficiency. In terms of generation realism, OneDC also achieves lower FID scores than all multi-step diffusion-based methods, ensuring visually pleasing, perceptually aligned reconstructions. On the MS-COCO 30K dataset, OneDC provides a 42.94% bitrate saving over PerCo—previously the highest-realism method—when evaluated by FID, further underscoring its effectiveness. In conclusion, OneDC consistently delivers faithful and perceptually compelling reconstructions across a variety of image resolutions and content complexities, despite relying on only a single sampling step. These results support our hypothesis that iterative sampling is not necessary for high-quality diffusion-based generative compression.
Qualitative Evaluation. Fig. 6 presents visual comparisons on the CLIC2020 test set. At the lowest bitrate, OneDC produces the most natural reconstructions with sharp high-frequency details, while MS-ILLM yields noticeably blurred textures. Multi-step diffusion-based methods (DiffEIC, PerCo, and DiffC) introduce structural distortions and deviated details. Similar trends are observed on the MS-COCO 30K dataset (Fig. 7), where MS-ILLM remains blurry and multi-step diffusion methods struggle to preserve visual fidelity. Additional examples are provided in the supplementary material.
Figure 7: Visual examples on the MS-COCO 30K dataset. Zoom in for better view.
[…]
Settings CLIC2020
DISTS FID
Semantic guidance
No guidance 44.0 45.1
Text guidance 24.2 36.3
Hyperprior guidance 20.7 24.3
Hyperprior + Sem. Distil.$\to$ Ours 0.00 0.00
Loss variation
Pixel-domain only 11.4 51.8
Latent-domain only 60.7 37.1
Hybrid-domain $\to$ Ours 0.00 0.00
[…]
Methods Times (s) MS-COCO 30K
Enc. Dec. LPIPS DISTS FID
VAE-based
MS-ILLM 0.14 0.17 76.9 225.3 247.9
Multi-step diffusion
DiffEIC 0.32 12.4 247.8 100.1 163.3
PerCo 0.58 8.80 410.8 265.9 75.2
DiffC 3.9$\sim$15.6 6.9$\sim$10.8 114.9 96.20 144.2
One-step diffusion
OneDC $\to$ Ours 0.15 0.34 0.00 0.00 0.00
Figure 8: Visual examples of the ablation study in hybrid-domain perceptual learning. No latent-domain training leads artifacts (Left), while no pixel-domain supervision leads color shift (Right).
4.3 Ablation Studies and Efficiency Evaluation
We conduct ablation studies to investigate the effectiveness of semantic guidance and hybrid-domain perceptual learning. Models are evaluated on the CLIC2020 test set with BD-Rate [5].
Semantic guidance. We first validate the importance of semantic guidance and then evaluate the effectiveness of our proposed semantic distillation for the hyperprior codec. As shown in Table 1, removing semantic guidance (“No Guidance”) significantly degrades reconstruction quality—by up to 40%—highlighting its critical role in our one-step diffusion model. While text-based guidance improves performance, it underperforms the hyperprior guidance on high-resolution CLIC2020 images. This suggests that the hyperprior’s locally captured semantics offer a more precise and spatially aligned representation of the image content. Further applying semantic distillation to the hyperprior (“Hyperprior + Sem. Distill.”) leads to additional gains, indicating that transferring prior knowledge from a pretrained tokenizer enhances the semantic capability of our hyperprior branch.
Hybrid-domain Perceptual Learning. We evaluate the effectiveness of hybrid pixel-latent domain training in Stage II, as summarized in Table 1. (1) Using the pixel-domain loss only (No $L_{\text{distill}}$ and $\gamma L_{adv}$) leads to notable drops in both FID and DISTS scores. Visual examples in Fig. 10-Left show grid-like artifacts, reflecting a significant decline in perceptual quality. This underscores the importance of diffusion distillation in enforcing distributional alignment for realistic reconstruction. (2) Using the latent-domain loss only (No $L_{\text{recon}}$) also results in consistent degradation across all metrics, accompanied by visible color shifts in the reconstructions (Fig. 10-Right). This highlights the importance of pixel-level supervision in image compression, which emphasizes accurate reconstruction—unlike the pure image generation task that focuses solely on visual realism.
Operation Efficiency. We evaluate the coding times of different methods on $1024\times 1024$ images using an A100 GPU. These runtime results are presented alongside BD-Rate results on the MS-COCO 30K dataset to offer a comprehensive view of operational efficiency. As shown in Table 2, OneDC achieves a significant decoding speed advantage—over 20×20$\times$20 × faster—compared to multi-step diffusion-based methods, while still maintaining impressive reconstruction quality. Although OneDC is slightly slower than the pure VAE-based MS-ILLM due to its larger model, this added capacity is crucial for preserving perceptual quality under extreme compression, as evidenced by its superior DISTS and FID performance. Overall, these results underscore the effectiveness of OneDC in balancing fidelity, realism, and efficiency at ultra-low bitrates.
5 Conclusion
We present OneDC, a novel one-step diffusion-based generative image codec designed for high-quality and efficient compression under ultra-low bitrate scenarios. OneDC integrates a latent compression module for efficient coding and a one-step diffusion generator for fast reconstruction. To provide the critical semantic guidance for the one-step diffusion model and address the limitations of textual guidance, we leverage the hyperprior as the substitute. Furthermore, we improve the semantic capability of the hyperprior through a distillation mechanism. To enhance training efficiency and stability, we adopt a two-stage strategy: Stage I focuses on compression training through pixel-domain supervision, while the Stage II refines reconstruction quality via hybrid-domain perceptual learning. Extensive experiments demonstrate that OneDC achieves SOTA perceptual quality with fast decoding, highlighting the strong potential of one-step diffusion models in generative image compression.
Limitation. Although OneDC offers substantial speedups over multi-step diffusion methods, its decoding speed does not meet real-time requirements. In future work, we plan to explore model distillation and architectural optimization to enhance efficiency.
Supplementary Material
Appendix A Experiment
Evaluation of third-party models. We evaluate MS-ILLM [39] using the official checkpoints and fine-tune them with the provided code to support lower bitrates. For GLC [18], we report the results directly from its paper because the code and models are not publicly available at present. For DiffEIC [28] and DiffC [49], we use the official implementations and released models. Specifically, we adopt the Stable Diffusion (SD) 2.1-based variant of DiffC, which shows slightly better performance compared to the SD 1.5 version. For PerCo [7], we use a publicly available reimplementation [23], as the original code has not been released. For multi-step diffusion codecs, we follow the default sampling settings provided in their code (e.g., 50 for DiffEIC, 20 for PerCo).
Test setting of FID metric. For the CLIC2020 test set (both full resolution and the $768\times 768$ cropped variant), as well as the DIV2K test set [1], we compute FID using $256\times 256$ patches, following the protocol in [36]. This setting is also consistent with that used in DiffEIC [28]. For the MS-COCO 30K dataset, we use $512\times 512$ patches for FID evaluation, following the setup in PerCo [7].
Additional datasets and metrics. To enable a more comprehensive comparison, we additionally evaluate our method on the DIV2K test set [1]. Notably, several diffusion-based codecs [28, 49] report results on the CLIC2020 test set [47] at a resolution of $768\times 768$ [28]. To ensure fair comparison, we include evaluations under this setting as well. Performance on DIV2K and CLIC2020 ($768\times 768$) is presented in Fig. 9, where OneDC consistently achieves SOTA results across datasets and metrics, demonstrating strong robustness and generalization.
We also report traditional pixel-level distortion metrics—PSNR and MS-SSIM [52]—to provide a more comprehensive analysis, as shown in Fig. 10. At extremely low bitrates, optimizing for PSNR often suppresses high-frequency details, resulting in blurred reconstructions [36]. While OneDC shows slightly lower PSNR than MS-ILLM, both perceptual metrics and qualitative examples clearly demonstrate its superior visual quality. Despite prioritizing perceptual realism, OneDC still achieves competitive PSNR compared to other diffusion-based methods. Moreover, on the MS-SSIM metric, OneDC matches MS-ILLM and outperforms all other baselines. These results confirm that OneDC delivers strong pixel-level fidelity alongside high perceptual realism, highlighting the overall effectiveness of our method.
Additional visual examples. We present more qualitative comparisons across four datasets: Kodak (Fig.13), CLIC2020—both full resolution (Figs.14) and $768\times 768$ cropped (Fig.15), and MS-COCO 30K (Fig.16). OneDC consistently outperforms prior SOTA methods, delivering superior visual quality across diverse content and resolutions, yet with the lowest bitrate cost.
Effectiveness of semantic distillation. To further evaluate the proposed semantic distillation strategy, we fine-tune the pretrained text-to-image one-step diffusion model [58], replacing its textual semantic condition with the hyperprior features produced by our semantic hyperprior decoder $h_{sem}$. This enables reconstruction using only the hyperprior signal. Fig. 17 presents additional reconstruction results on the COCO2017 validation dataset [30], extending Fig. 2 from the main paper. These results confirm that the distilled model captures richer semantic information, facilitating final reconstruction.
Figure 9: Comparison of methods on the CLIC2020 test ($768\times 768$) and the DIV2K dataset.
Figure 10: Comparison of methods measured by PSNR and MS-SSIM.
Appendix B Training Details
Stage I training. This stage focuses on training the compression module and fine-tuning the one-step diffusion model [58] for the image reconstruction task. The training loss is defined as:
$L_{stageI}=L_{recon}+\lambda R+\alpha L_{aux},\;\;\;\;\text{where}\;\;L_{recon}=L_{1}+L_{perceptual}$ (5)
We use the L1 as the pixel-level loss and the LPIPS [19] as the perceptual-level loss. To support various bitrates, the rate-distortion trade-off parameter $\lambda$ is set to $\{1.8,2.9,4.6,7.4\}$. An auxiliary code prediction loss $L_{aux}$ is included with a weighting factor of $\alpha=0.001$. We train our model on the dataset introduced in [14]. Training is performed on 4×A100 GPUs for 800,000 steps, using a three-stage learning rate schedule with AdamW [33]: a) 5e-5 for the first 500,000 steps; b) 1e-5 for the next 200,000 steps; c) 1e-6 for the final 100,000 steps. During training, image patches of size $\{512,1024\}$ are randomly cropped with probabilities of $\{0.6,0.4\}$, respectively. The batch size is set to 32 for $512\times 512$ crops and 8 for $1024\times 1024$ crops (across 4 GPUs).
Stage II training. This stage fine-tunes the one-step generator to improve reconstruction realism by better aligning the distribution of its outputs with that of real images. The training objective is:
$L_{stageII}=L_{distill}+\beta L_{recon}+\gamma L_{adv},\;\;\;\;\text{where:}$: (6)
$L_{distill}=\mathbb{E}_{t,\tilde{y}^{\prime}_{t}}[\epsilon_{fake}(\tilde{y}^{\prime}_{t},t)-\epsilon_{real}(\tilde{y}^{\prime}_{t},t)],\;\;L_{adv}=\mathbb{E}_{t,\tilde{y}^{\prime}_{t}}[-Disc(\epsilon_{fake}(\tilde{y}^{\prime}_{t},t),t)]$ (7)
Here, $L_{distill}$ represents the diffusion distillation loss [59], and $L_{adv}$ is the adversarial loss in the latent space, following [58]. $Disc$ denotes the discriminator network, which takes the mid feature in the diffusion U-Net as the input [58]. The variable $\tilde{y}_{0}$ is the latent output generated by the one-step diffusion model, and $\tilde{y}^{\prime}_{t}$ is its noised version at timestep $t$. We uniformly sample $t\in[20,640]$, since synthesizing high-frequency details does not require large noise levels. The weighting parameters are set as follows: $\beta=0.625$ balance the reconstruction and distillation terms, and $\gamma=0.001$ for the adversarial loss ($\gamma$ follows [58]).
Specifically, diffusion distillation [59] minimizes the expected Kullback-Leibler (KL) divergence between the time-dependent distributions of the target $p_{real,t}$ and the generator output $p_{real,t}$, thereby effectively transferring knowledge from the multi-step diffusion model to the one-step generator. The gradient used to update the one-step generator parameters $\theta$ is given by the difference between the score functions of the real and fake distributions:
$\nabla_{\theta}L_{distill}$ $=\mathbb{E}_{t,\tilde{y}_{0}}(\nabla_{\theta}\text{KL}(p_{fake,t}||p_{real,t}))$ (8)
$=-\mathbb{E}_{t,\tilde{y}^{\prime}_{t}}[(s_{real}(\tilde{y}^{\prime}_{t},t)-s_{fake}(\tilde{y}^{\prime}_{t},t))\frac{d\epsilon_{\theta}}{d\theta}]$ (9)
$=\mathbb{E}_{t,\tilde{y}^{\prime}_{t}}[(\epsilon_{fake}(\tilde{y}^{\prime}_{t},t)-\epsilon_{real}(\tilde{y}^{\prime}_{t},t))\frac{d\epsilon_{\theta}}{d\theta}]$ (10)
Here, $s_{\text{real}}$ and $s_{\text{fake}}$ are the score functions learned by multi-step diffusion model $\epsilon_{\text{real}}$ and $\epsilon_{\text{fake}}$ respectively. To ensure the fake score network $\epsilon_{\text{fake}}$ accurately tracks the evolving distribution of the one-step diffusion model, we update it using a standard denoising loss:
$L_{fake}=\mathbb{E}_{t,\tilde{y}^{\prime}_{t}}||\epsilon_{fake}(\tilde{y}^{\prime}_{t},t)-\tilde{y}_{0}||^{2}_{2}$ (11)
The improved version of diffusion distillation proposed in [58] introduces adversarial training in the latent space to further enhance distribution alignment. A discriminator is trained to differentiate between features extracted from real and generated images, using the following objective:
$L_{Disc}=\mathbb{E}_{t,x}[Disc(\epsilon_{fake}(\tilde{y}^{\prime}_{t},t))-Disc(\epsilon_{fake}(E_{\text{VAE}}(x)+n_{t},t))]$ (12)
Here, $E_{\text{VAE}}$ denotes the encoder of the diffusion model, and $n_{t}$ is the sampled noise in timestep $t$. The fake branch is then optimized with an additional adversarial loss:
$L=L_{fake}+\sigma L_{Disc}$ (13)
We follow [58] and set $\sigma=0.01$. During training, the fake branch and discriminator are each updated 10 times for every update of the one-step generator, ensuring stable adversarial optimization.
Training is conducted on 4×A100 GPUs for 1,000,000 steps. The learning rate is fixed at 1e-6 (with AdamW) for the one-step generator, fake network, and discriminator. The batch size, cropping strategy and training data are identical to those used in Stage I.
Appendix C Model Details
The overall architecture is shown in Fig.11. To ensure better alignment with the latent space of the diffusion model, we extract features from its pretrained VAE encoder. Additionally, inspired by DiffEIC[28], we incorporate embeddings from the original input image to enrich the encoder with complementary spatial and semantic information. To jointly learn compressible latents and capture high-level semantics, we introduce a lightweight U-Net within the analysis transform. Its multi-scale design facilitates effective aggregation of both local textures and global structures. The extracted semantic features are integrated into the hyperprior branch, enhancing its representational capability.
For parameter-efficient adaptation, we insert LoRA [17] layers across all modules of the pretrained one-step diffusion U-Net, setting the LoRA rank to 64. The combined parameter count of the encoder and decoder components ($g_{a},h_{enc},h_{ctx},h_{sem}$ and $g_{s}$) is 394M, while the adapted diffusion U-Net contributes an additional 928M parameters.
Further details of the semantic distillation mechanism are provided in Fig. 12. We adopt a Swin Transformer [32] to improve adaptability across different image resolutions during training.
Figure 11: Details of our model architecture. The pratrained VAE encoder is from SD 1.5. For the U-Net used in $g_{a}$, we use the implementation from the diffusers library [48].
Figure 12: Details of the auxiliary code predictor module. For Swin Transformer block [32], we set window size to 16, head dim to 64. W means normal window, SW means shifted window.
Appendix D Social Impact
Positive Aspects. Generative codecs offer substantial benefits by significantly reducing the storage and transmission demands of high-resolution imagery through the synthesis of perceptually convincing content. This improvement in efficiency can help broaden access to high-quality visual media, particularly in bandwidth-limited or resource-constrained environments. The resulting gains in data economy and speed of delivery have promising implications for applications in social communication and personal entertainment.
Negative Aspects. Despite these advantages, generative compression introduces synthesized content that may deviate from the original input, raising concerns about the fidelity and authenticity of reconstructed images. Ongoing research into multi-realism codecs [3] offers a potential path forward, enabling more transparent control over the trade-off between realism and fidelity.
Figure 13: Visual results on the Kodak dataset. The VAE-based MS-ILLM exhibits noticeable artifacts. Compared with our OneDC, the previous SOTA multi-step diffusion codec DiffC requires at least 1.4×1.4$\times$1.4 × bitrate on these two examples. Zoom in for better view.
Figure 14: Visual results on the CLIC2020 test set. Zoom in for better view.
Figure 15: Visual results on the CLIC2020 test set ($768\times 768$). Compared with OneDC, the previous SOTA multi-step diffusion method DiffC requires $\geq 1.5$ higher bitrate while still producing slightly lower reconstruction quality on these examples. Zoom in for better view.
Figure 16: Visual results on the MS-COCO 30K dataset. PerCo, the previous best-performing method in terms of FID, requires $\geq 1.4$ bitrate over OneDC, but still results in suboptimal fidelity. Zoom in for better view.
Figure 17: Reconstructions from different semantic guidance. (a) Text prompts struggle to capture complex visual semantics, resulting in severe distortions when using a pretrained text-to-image one-step diffusion model [58]. (b) We finetune the model [58] for hyperprior-to-image generation. Hyperprior guidance yields more faithful reconstructions. (c) Our proposed semantic distillation further improves object-level accuracy.