### v2

# One-Step Diffusion-Based Image Compression with Semantic Distillation

Abstract
[…]
footnotetext: Naifu Xue and Zhaoyang Jia are visiting students at Microsoft Research Asia.
1 Introduction
The rapid growth in image data has led to increased storage and transmission costs, heightening the need for efficient, high-quality image compression techniques. While VAE-based learned image compression (LIC) [36, 71] has surpassed traditional codecs (e.g., VVC [8]) in rate-distortion (RD) performance, it often produces blurry details at low bitrates [41] since they are typically optimized for objective distortion rather than human perception [7]. To mitigate this, some approaches [41, 44] focus on optimizing visual quality by incorporating perceptual metrics and adversarial losses to enable generative image compression. Despite these advances, such generative models tend to introduce artifacts under extreme compression, degrading the realism of reconstructed images.
Recently, diffusion-based generative codecs [9, 33] have been introduced to enhance reconstruction quality by leveraging the powerful content synthesis capabilities of pretrained models. While these methods significantly improve perceptual realism, they may occasionally generate content that deviates from the original input, thereby compromising reconstruction fidelity. In addition, their inherently iterative sampling process leads to substantial computational overhead, making them notably slower than conventional VAE-based codecs.
In standard diffusion-based image generation, the model progressively refines a noisy signal through iterative denoising, beginning with coarse structures and gradually synthesizing high-frequency details [65]. However, when applied to image compression, the task changes: at low bitrates, codecs typically retain coarse structures in the compressed latents. As a result, the decoder is primarily responsible for plausibly reconstructing high-frequency content from the preserved low-frequency information. This observation motivates our central hypothesis: given the compressed latent, multi-step sampling is not necessary for decoding, and a carefully designed one-step alternative could suffice. Although recent advances in one-step diffusion [53, 66, 25] offer promising tools for accelerating inference, their potential remains underexplored in the context of image compression.
Figure 1: Top: multi-step sampling is not essential for image compression; intermediate results are from DiffEIC [33]. Bottom: Visual comparisons including existing open-sourced multi-step diffusion codecs [9, 33, 57] and our proposed one-step codec. Our method achieves the highest visual quality at the lowest bitrate while offering significantly faster decoding.
In this paper, we introduce OneDC (One-step Diffusion-based generative Codec), a novel framework for ultra-low bitrate image coding. OneDC integrates a latent compression module with a one-step diffusion generator: the former encodes the image into compact latents, while the latter synthesizes high-frequency details conditioned on the latent. Since the pixel-domain training is insufficient for guaranteeing perceptual quality at low bitrates [22], we adopt a hybrid-domain training strategy. Specifically, we combine a pixel-domain perceptual loss to promote fidelity with a latent-domain diffusion distillation objective [66] to enhance realism. This design allows OneDC to effectively leverage the pretrained diffusion model while achieving a balanced trade-off among compression ratio, perceptual quality, and decoding efficiency, as illustrated in Fig. 1.
Moreover, we further explore the role of semantic guidance (i.e., the input of the cross-attention layers [50]) within our framework. In one-step diffusion, such guidance is essential to compensate for the absence of multi-step refinement (see ablation in Section 4.3). While existing diffusion models have utilized text prompts for conditioning, we argue that text is suboptimal for image compression due to two limitations: (1) natural language struggles to capture fine-grained or localized visual semantics, and (2) generating high-quality captions typically requires large-scale vision-language models (e.g., the large BLIP2 [32] used in [9]), introducing substantial computational overhead. This raises a central question: Can we design more effective semantic representations than text for guiding one-step diffusion-based codecs—without incurring excessive computational cost?
Recent studies have shown that the hyperprior in VAE-based codecs can capture high-level semantic information [22, 48]. Compared to textual prompts, hyperpriors provide more precise and spatially aligned semantic cues, making them well-suited for representing localized semantics in high-resolution images (see Fig.2). This insight motivates us to further enhance the representational capability of the hyperprior. Inspired by generative tokenizers [15], where codebooks encode rich semantic content, we propose hyperprior semantic distillation—a training strategy that transfers semantic knowledge from a pretrained tokenizer to the hyperprior through a code prediction module [72]. As illustrated in Fig.2, this approach results in reconstructions that are more semantically accurate and visually coherent.
[…]
Figure 2: Reconstructions from different semantic guidance. (a) Text prompts (from GPT-4o [47]) struggle to capture complex visual semantics, resulting in severe distortions when using a pretrained text-to-image one-step diffusion model [66]. (b) We finetune the model [66] for hyperprior-to-image generation. Hyperprior guidance yields more faithful reconstructions. (c) Our proposed semantic distillation further improves object-level accuracy, particularly in the highlighted regions.
  • We propose OneDC, a one-step diffusion codec comprising a latent compression module for compact feature encoding and a one-step diffusion generator for fast, high-quality decoding. A hybrid-domain training strategy further enhances both fidelity and perceptual realism.
  • We identify the importance of high-level semantic guidance in one-step diffusion and highlight the limitations of text embeddings. To improve reconstruction quality, we introduce hyperprior features as an alternative and enhance them through semantic distillation.
  • Extensive experiments show that OneDC achieves SOTA compression performance while offering significantly faster decoding than existing diffusion-based codecs, demonstrating the potential of one-step diffusion in generative compression.
2 Related Work
VAE-based Learned Image Compression. Learned image compression has made rapid progress in RD performance. Ballé et al. [4] introduced an end-to-end framework with analysis/synthesis transforms and a factorized entropy model. The hyperprior model [5] enhanced entropy modeling by encoding side information (hyperprior) to predict distribution parameters, substantially improving RD performance. Subsequent work combined hyperpriors with spatial context [43] and adopts more expressive distribution models such as Gaussian mixtures [11]. Recent studies further refined transform architectures for stronger nonlinear representation [36] and developed advanced entropy models for more compact symbol coding [23], together surpassing the VVC-Intra baseline [8]. Nevertheless, these models are typically optimized with distortion-oriented objectives such as MSE or MS-SSIM [60], which often over-smooth textures and degrade perceptual quality at low bitrates [41], even with SOTA transforms and entropy models [36, 71, 58].
VAE-based Generative Image Compression. To improve perceptual quality at low bitrates, GAN-based approaches such as Agustsson et al. [2] and HiFiC [41] train VAE codecs to synthesize high-frequency details using perceptual losses [24] and adversarial discriminators. MS-ILLM [44] improves realism through a local likelihood model, while EGIC [27] introduces semantic-aware feedback via segmentation-conditioned discriminators. Other methods enhance compression by integrating generative tokenizers: Mao et al. [40] utilize VQGAN [15] for compression, Xue et al. [63] unify token generation and entropy coding, and Jia et al. [22] propose GLC that improves token compression via transform coding. More recently, dual-branch codecs [39, 64] have been introduced to support more compact and flexible latent representations. While these methods highlight the strong feature learning capacity of VAE-based codecs, their reconstruction quality at extremely low bitrates remains limited by model capacity and training scale.
Diffusion-based Generative Image Compression. Diffusion models have rapidly evolved in recent years, progressing from theoretical foundations [51, 20] to practical high-quality implementations [50]. Building on these advances, recent studies [29, 9, 33, 57, 62] have explored the use of generative priors from large-scale pretrained diffusion models in image compression, demonstrating improved perceptual realism compared to conventional VAE-based approaches. Eric et al. [29] combine pretrained diffusion with ControlNet [69] to reconstruct images from captions and compressed sketches. PerCo [9] fine-tunes diffusion models using captions and quantized features, with captions generated by a large BLIP2 model [32]. DiffEIC [33] conditions diffusion models on VAE latents, showing strong performance even without text, and their subsequent work [34] improves encoder and introduces relay fine-tuning. DiffC [57] demonstrates zero-shot compression with diffusion models using reverse-channel coding. DDCM [46] replaces the continuous Gaussian noise in sampling with discrete codebook selections for compression task. While these approaches improve realism at ultra-low bitrates, they often struggle to maintain high fidelity. Moreover, multi-step sampling in DDIM [52] introduces substantial decoding overhead, limiting their practicality. By contrast, our OneDC achieves high fidelity and realism with significantly faster decoding via one-step sampling. While prior work [33] suggests high-level semantics (e.g., text) may be optional, we find such guidance remains crucial in our one-step diffusion for image decoding task, as detailed in Section 4.3.
One-step Diffusion Models. To reduce the computational cost, recent efforts have focused on distilling multi-step models into efficient one-step generators. Song et al. [53] introduce Consistency Models, which achieve few-step generation by enforcing self-consistency along diffusion trajectories. Yin et al. [67] propose Distribution Matching Distillation (DMD), which minimizes an approximate KL divergence between real and generated image distributions modeled by multi-step teacher networks. Their subsequent work, DMD2 [66], further improves generation quality by introducing a diffusion-GAN framework and enhanced training strategies. Based on DMD, Song et al. [54] introduce class-specific student models to improve performance across diverse categories. SwiftBrush series [45, 12] adopt a LoRA adaptation [21] for variational score distillation and further optimize diversity by integrating a clamped CLIP loss. Independently, Kang et al. [25] employ perceptual loss supervision and a multi-scale U-Net discriminator for one-step distillation. These models have also been extended to low-level vision tasks such as super-resolution [14, 61, 59, 26] and image restoration [18], highlighting the potential of one-step diffusion beyond generation. Inspired by these advances, we explore its application in generative image compression, a setting that remains underexplored.
3 Methodology
Figure 3: Overview of the OneDC framework. Q denotes scalar quantization, and FSQ stands for finite scalar quantization. AE and AD refer to the arithmetic encoder and decoder, respectively. $h_{ctx}$ and $h_{sem}$ represent the context and semantic decoders used in the hyperprior branch.
We propose OneDC, a one-step diffusion-based generative image compression framework that achieves high-quality reconstruction with low-latency decoding under ultra-low bitrate constraints. Section 3.1 introduces the coding pipeline of our method, consisting of a latent compression module that encodes the input image into a compact latent representation and a one-step diffusion generator that synthesizes reconstructions with realistic details. Given the critical role of semantic guidance in our one-step diffusion and the limitations of textual prompts in compression scenarios, Section 3.2 introduces the use of the hyperprior as an alternative conditioning signal. To further tap the potential of the hyperprior, we propose a semantic distillation strategy that transfers knowledge from a pretrained generative tokenizer to the hyperprior codec, thereby improving semantic accuracy. Finally, Section 3.3 presents our training scheme, which combines pixel-domain and latent-domain objectives to jointly optimize reconstruction fidelity and perceptual realism.
3.1 Framework Overview
[…]
Figure 4: Two stage training pipeline of OneDC. The codebook in semantic distillation is initialized from the pretrained tokenizer, and the discriminator in diffusion distillation is abbreviated as Disc.
[…]
3.2 Semantic Guidance with Hyperprior
3.2.1 From Text to Hyperprior
In the one-step diffusion setting, semantic guidance (the input of the cross-attention layers [50]) plays an even more critical role than in conventional multi-step diffusion. Unlike multi-step methods, which iteratively refine the output and can gradually correct semantic inconsistencies, one-step diffusion models rely entirely on a single forward pass, making accurate semantic conditioning essential. While existing diffusion models typically rely on textual prompts to provide high-level semantic guidance, such prompts are suboptimal in the context of natural image compression. Text struggles to capture localized semantics and often fails to describe fine-grained visual attributes—such as object boundaries and textures—particularly in high-resolution imagery. In addition, generating text descriptions typically requires large models [32], introducing significant computational overhead.
Figure 5: Visual examples on the CLIC2020 test set. Zoom in for better view.
Recent work [22] suggests that the categorical distributed hyperprior with vector quantization (VQ) can capture semantic information. Building on this insight, OneDC replaces textual prompts with the categorical hyperprior $\hat{z}$ as semantic guidance for the diffusion model. To compress the hyperprior, we adopt finite scalar quantization (FSQ) [42], a simpler yet more effective alternative to VQ [15], to learn a categorical $\hat{z}$ with semantic information. The hyperprior $z$ is compressed into 7 channels with 4 quantization levels each, yielding an equivalent codebook size of 16,384. At a $64\times$ spatial downsampling, this design results in a bitrate of just 0.0034 bpp, incurring very low transmission overhead. To substitute for text embeddings, we introduce a semantic decoder $h_{sem}$ that transforms the quantized hyperprior $\hat{z}$ into a semantic context $c\in\mathbb{R}^{B\times N\times D}$, where $N=H^{\prime}\times W^{\prime}$ is the flattened spatial shape and $D$ is the embedding dimension. This context is injected into the cross-attention layers of the one-step diffusion U-Net, replacing the textual embeddings used in the original design. In each cross-attention layer, the input latent feature $f_{in}$ serves as the query, while the semantic context $c$ provides both keys and values:
$f_{out}=\mathrm{Softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V,\quad\text{where}~Q=W_{Q}f_{in},~K=W_{k}c,~V=W_{v}c$ (1)
where $d_{k}$ denotes the dimensionality of the key vectors and $f_{out}$ is the output feature of the cross-attention layer. This design enables every spatial location in the latent to attend adaptively to semantically relevant tokens, strengthening content alignment. Compared with text conditioning, hyperprior-based guidance offers three advantages: (1) As shown in Fig. 2, conditioning the one-step diffusion model on categorical hyperprior tokens yields more faithful reconstructions than using text prompts, validating their effectiveness as semantic guidance. (2) The $64\times$ downsampled hyperprior provides a large receptive field while retaining spatial locality, thereby offering a more balanced global–local semantics compared to purely global text embeddings. (3) Unlike separately trained text encoders [49], our approach enables end-to-end optimization of semantic guidance, allowing the hyperprior to adapt jointly with the diffusion model for improved content expression.
3.2.2 Semantic Distillation for Hyperprior
While the hyperprior effectively captures semantic cues for diffusion guidance, its representation capability can be further improved. In particular, we propose a semantic distillation mechanism aimed at fully tapping the potential of the hyperprior branch. This distillation guides the hyperprior encoder to better capture visual content and improve the decoding quality of the semantic decoder $h_{sem}$. Specifically, we transfer knowledge from a pretrained high-capacity generative tokenizer [15] to the hyperprior codec via an auxiliary distillation task. This choice is motivated by the structural similarity between the categorical hyperprior codec and the generative tokenizers that are known to produce semantically rich discrete representations [68]. By leveraging this compatibility, we use the tokenizer as a teacher to guide the hyperprior toward more effective semantic encoding.
Following GLC [22], we introduce a transformer-based predictor $P_{aux}$ [72], trained to predict discrete token labels from the hyperprior-derived semantics $c$. The ground truth labels $I_{gt}$ are obtained from a pretrained tokenizer encoder $E_{aux}$. Since the tokenizer is optimized for perceptual reconstruction, its objectives align well with OneDC, making it an effective teacher. Moreover, the smaller information bottleneck in the hyperprior branch naturally filters out redundant information, distilling the most salient semantic features from the teacher. The distillation is supervised via cross-entropy (CE) loss:
$I_{gt}=VQ(E_{aux}(x)),\;\;\;\;L_{aux}=CE(I_{gt},P_{aux}(c))$ (2)
Both $P_{aux}$ and $E_{aux}$ are used only during training, introducing no inference overhead. As shown in Fig.2, our semantic distillation enhances content fidelity, and ablation results in Section4.3 further validate its effectiveness in improving final reconstruction quality.
Figure 6: Visual examples on the MS-COCO 30K dataset. Zoom in for better view.
3.3 Training Strategy
To enhance training efficiency and stability, we adopt a two-stage training strategy, inspired by prior works [41, 44]. As shown in Fig. 4, stage I mainly trains the compression module while stage II focuses on fine-tuning the one-step diffusion model for better reconstruction quality. In both stages, the one-step diffusion model is adapted via LoRA layers [21], allowing fast convergence for the compression task while preserving rich generative priors.
Stage I: Pixel-domain Compression Learning This stage aims to: (1) train the codec to compress images into compact latent representations with high-fidelity reconstruction, (2) embed semantic information into the hyperprior via distillation, and (3) adapt the one-step diffusion model to synthesize fine-grained details on the decoder side. The total loss is defined as:
[…]
We use LPIPS [70] as the perceptual loss and $L_{1}$ as the pixel loss. $R$ is the bitrate loss from the quad-tree-based spatial entropy model [30, 31], and $L_{aux}$ is the semantic distillation loss (Eq. 2). The hyperparameter $\lambda$ controls the compression ratio and $\alpha$ is the weight of semantic distillation loss.
Figure 7: Rate–distortion curves under the full-resolution setting. Zoom in for better view.
Figure 8: Rate–distortion curves under the resize & center-crop on Kodak $\times512$ and CLIC2020 $\times768$. Zoom in for better view.
Stage II: Hybrid-domain Perceptual Learning. This stage fine-tunes the one-step diffusion model to enhance perceptual realism without degrading fidelity. The latent compression module is fixed to improve training stability [41]. Prior work has shown that pixel-level optimization alone is insufficient for optimal perceptual quality [22]. Thus, we adopt a joint training strategy in both pixel and latent domains. Specifically, we integrate the diffusion-based distillation method from [66], which transfers generation knowledge from a pretrained multi-step diffusion teacher. The training also involves both perceptual supervision from the pixel domain and adversarial alignment in the latent domain, which ensures fidelity by aligning reconstruction with the original image. The total loss is:
$L_{stageII}=L_{distill}+\beta L_{recon}+\gamma L_{adv},\;\;\;\;\text{where:}$ (4)
$L_{distill}=\mathbb{E}_{t,\tilde{y}^{\prime}_{t}}[\epsilon_{fake}(\tilde{y}^{\prime}_{t},t)-\epsilon_{real}(\tilde{y}^{\prime}_{t},t)],\;\;L_{adv}=\mathbb{E}_{t,\tilde{y}^{\prime}_{t}}[-Disc(\epsilon_{fake}(\tilde{y}^{\prime}_{t},t),t)]$ (5)
Here, $\epsilon_{real}$ and $\epsilon_{fake}$ denote the real and fake score networks in the teacher model, and $\tilde{y}^{\prime}_{t}$ is the noised latent at timestep $t$. DiscDisc refers to the discriminator. $\beta$ and $\gamma$ balance the pixel and adversarial losses. Additional implementation and loss details are provided in the supplementary material.
4 Experiment
4.1 Implementation and Evaluation Settings
Model and Training. Our generator adopts the U-Net architecture from the Stable Diffusion 1.5 (SD1.5) [50] and is initialized with weights from a pretrained one-step text-to-image task [66]. In Stage I, we use the pretrained tokenizer from MaskGIT [10] for semantic distillation. In Stage II, we employ the multi-step SD1.5 pretrained model as the teacher for diffusion distillation. To improve high-resolution adaptability, we randomly crop patches of size 512 or 1024 during training. Models are optimized using AdamW [38]. Additional settings are provided in the supplementary material.
Compared methods. We compare our OneDC with generative image codecs at a low-bitrate scenario. The compared methods include the vanilla VAE-based generative codec MS-ILLM [44] and GLC [22], and the recent diffusion-based methods: PerCo (SD) [28], DiffEIC [33], DiffC [57] (SD2.1-based) and DDCM [46]. For most methods, we evaluate at full-resolution, consistent with previous implementations in generative neural codecs [41, 44, 22]. For some diffusion-based methods like DDCM, we additionally provide results under the resize & center-crop setting, following DDCM protocol [46] for fair comparison. Further details are provided in the supplementary material.
Evaluation datasets and metrics. We evaluate OneDC on several datasets, including Kodak [16], CLIC2020 test set [55], and MS-COCO 30K [35]. Reconstruction fidelity is evaluated using perceptual metrics LPIPS [24] and DISTS [13], along with the traditional metrics PSNR and MS-SSIM [60], while generative realism is measured by the no-reference perceptual metric FID [19]. Bitrate saving is measured by BD-Rate [6]. It is worth noting that FID is computed on overlapping $256\times 256$ patches at full-resolution CLIC2020, following previous practice [44, 22]. For the MS-COCO 30K dataset with $512\times 512$ images, we evaluate FID on full-images, consistent with [9, 28]. However, large patch leads to unreliable FID on the small test set (Kodak with 24 images) [9]. To address this, under the resize & center-crop setting we follow the DDCM  [46] and compute FID on overlapping $64\times 64$ patches for Kodak. More details are provided in the supplementary material.
4.2 Main Results
Quantitative Evaluation. Fig. 7 presents the full-resolution comparison between OneDC and existing methods across multiple distortion metrics. Across all datasets, including high-resolution CLIC2020, low-resolution Kodak and MS-COCO 30K, OneDC consistently achieves the best reference-based perceptual fidelity, as measured by LPIPS and DISTS. Compared with the previous SOTA, multi-step diffusion-based DiffC, OneDC achieves 55.27% bitrate saving on Kodak and 54.60% on CLIC2020 in terms of LPIPS, demonstrating superior compression efficiency. On traditional metrics such as PSNR, OneDC also exhibits competitive performance. For generation realism, OneDC attains lower FID than all multi-step diffusion baselines when FID is evaluated on the CLIC2020 test set. On the MS-COCO 30K dataset, OneDC provides a 39.55% bitrate saving over PerCo (SD), previously the highest-realism method on this dataset. Such result further underscoring our method’s effectiveness. Furthermore, Fig. 8 reports results under the resize & center-crop setting: OneDC achieves better LPIPS and PSNR than DDCM while maintaining competitive realism, demonstrating strong fidelity-oriented perceptual compression. In conclusion, OneDC delivers faithful and perceptually compelling reconstructions across a wide range of resolutions and content complexities, despite relying on a single sampling step. These evaluation results support our hypothesis that iterative sampling is unnecessary for high-quality diffusion-based generative compression.
Qualitative Evaluation. Fig. 5 presents visual comparisons on the CLIC2020 test set (full-resolution). At the lowest bitrate, OneDC produces the most natural and faithful reconstructions with sharp high-frequency details, while MS-ILLM yields noticeably blurred textures and multi-step diffusion-based methods (DiffEIC, PerCo, and DiffC) introduce structural distortions and inconsistent details. Similar trends are observed on the MS-COCO 30K dataset (Fig. 6), where MS-ILLM remains blurry and multi-step diffusion methods struggle to preserve visual fidelity. These qualitative results, together with the quantitative comparisons, highlight the SOTA performance of the proposed OneDC. Additional visual examples are provided in the supplementary material.
Table 1: Ablation studies with BD-Rate ($\%$) $\downarrow$.
Settings CLIC2020
DISTS FID
Semantic guidance
No guidance 44.0 45.1
Text guidance 24.2 36.3
Hyperprior guidance 20.7 24.3
[…]
Loss variation
Pixel-domain only 11.4 51.8
Latent-domain only 60.7 37.1
Hybrid-domain $\to$ Ours 0.00 0.00
Table 2: Comparison of coding time and BD-Rate ($\%$) $\downarrow$.
Methods Times (s) MS-COCO 30K
Enc. Dec. LPIPS DISTS FID
VAE-based
MS-ILLM 0.14 0.17 138.3 253.0 478.4
Multi-step diffusion
DiffEIC 0.32 12.4 305.0 239.1 341.0
PerCo (SD) 0.58 8.80 538.8 345.8 59.6
[…]
One-step diffusion
OneDC $\to$ Ours 0.15 0.34 0.00 0.00 0.00
Notes: Metrics in the tables denote those used for BD-Rate calculation [6]; Ours is the anchor (0.00%).
Figure 9: Visual examples of the ablation study in hybrid-domain perceptual learning. No latent-domain training leads artifacts (Left), while no pixel-domain supervision leads color shift (Right).
4.3 Ablation Studies and Efficiency Evaluation
We conduct ablation studies to investigate the effectiveness of semantic guidance and hybrid-domain perceptual learning. Models are evaluated on the CLIC2020 test set with BD-Rate [6].
Semantic guidance. We first validate the importance of semantic guidance and then evaluate the effectiveness of our proposed semantic distillation for the hyperprior codec. As shown in Table 1, removing semantic guidance (“No Guidance”) significantly degrades reconstruction quality—by up to 40%—highlighting its critical role in our one-step diffusion model. While text-based guidance improves performance, it underperforms the hyperprior guidance on high-resolution CLIC2020 images. This suggests that the hyperprior’s locally captured semantics offer a more precise and spatially aligned representation of the image content. Further applying semantic distillation to the hyperprior (“Hyperprior + Sem. Distill.”) leads to additional gains, indicating that transferring prior knowledge from a pretrained tokenizer enhances the semantic capability of our hyperprior branch.
Hybrid-domain Perceptual Learning. We evaluate the effectiveness of hybrid pixel-latent domain training in Stage II, as summarized in Table 1. (1) Using the pixel-domain loss only (No $L_{distill}$ and $L_{adv}$) leads to notable drops in both FID and DISTS scores. Visual examples in Fig. 9-Left show grid-like artifacts, reflecting a significant decline in perceptual quality. This underscores the importance of diffusion distillation in enforcing distributional alignment for realistic reconstruction. (2) Using the latent-domain loss only (No $L_{recon}$) also results in consistent degradation across all metrics, accompanied by visible color shifts in the reconstructions (Fig. 9-Right). This highlights the importance of pixel-level supervision in image compression, which emphasizes accurate reconstruction—unlike the pure image generation task that focuses solely on visual realism.
Operation Efficiency. We evaluate the coding times of different methods on $1024\times 1024$ images using an A100 GPU. These runtime results are presented alongside BD-Rate results on the MS-COCO 30K dataset to offer a comprehensive view of operational efficiency. As shown in Table 2, OneDC achieves a significant decoding speed advantage—over $20\times$ faster—compared to multi-step diffusion-based methods, while still maintaining impressive reconstruction quality. Although OneDC is slightly slower than the pure VAE-based MS-ILLM due to its larger model, this added capacity is crucial for preserving perceptual quality under extreme compression, as evidenced by its superior DISTS and FID performance. Overall, these results underscore the effectiveness of OneDC in balancing fidelity, realism, and efficiency at ultra-low bitrates.
5 Conclusion
We present OneDC, a novel one-step diffusion-based generative image codec designed for high-quality and efficient compression under ultra-low bitrate scenarios. OneDC integrates a latent compression module for efficient coding and a one-step diffusion generator for fast reconstruction. To provide the critical semantic guidance for the one-step diffusion model and address the limitations of textual guidance, we leverage the hyperprior as a substitute. Furthermore, we improve the semantic capability of the hyperprior through a distillation mechanism. To enhance training efficiency and stability, we adopt a two-stage strategy: Stage I focuses on compression training through pixel-domain supervision, while Stage II refines reconstruction quality via hybrid-domain perceptual learning. Extensive experiments demonstrate that OneDC achieves SOTA perceptual quality with fast decoding, highlighting the strong potential of one-step diffusion models in generative image compression.
Limitation. Although OneDC offers substantial speedups over multi-step diffusion methods, its decoding speed does not meet real-time requirements. In future work, we plan to explore model distillation and architectural optimization to enhance efficiency.
Appendix A Experiment
Evaluation of third-party models. We evaluate MS-ILLM [44] using the official checkpoints and fine-tune them with the provided code to support lower bitrates. For GLC [22], we report the results directly from its paper because the code and models are not publicly available at present. For DiffEIC [33] and DiffC [57], we use the official implementations and released models. Specifically, we adopt the Stable Diffusion (SD) 2.1-based variant of DiffC, which shows slightly better performance compared to the SD 1.5 version. For PerCo [9], we use a publicly available reimplementation PerCo (SD) [28], as the original code has not been released. We also report the comparison with original PerCo by extract data from their paper (Fig. 10). For DDCM [46], we use the number reported in their paper. For multi-step diffusion codecs, we follow the default sampling settings provided in their code (e.g., 50 for DiffEIC, 20 for PerCo).
Test settings and FID calculation. At the full-resolution setting, we compute FID using overlapping $256\times 256$ patches for the CLIC2020 and DIV2K test sets [1], following the protocol of [41]. For the MS-COCO 30K dataset with $512\times 512$ images, FID is evaluated on entire images, consistent with [9, 28]. At the resize & center-crop setting, we resize the short side of each image (512 for Kodak, 768 for CLIC2020 test set) and then apply a center crop. In this setting, we use $64\times 64$ patches for FID calculation on Kodak and $128\times 128$ patches on CLIC2020 test set, consistent with DDCM [46]. It is worth noting that DiffEIC also evaluates under the resize & center-crop setting on CLIC2020 test set at 768 resolution, but computes FID with overlapping $256\times 256$ patches. For completeness, we additionally report results under DiffEIC’s protocol, denoted as resize & center-crop, 256 FID.
Additional datasets and metrics. To enable comprehensive comparison, we further evaluate our method on the DIV2K test set [1] under the full-resolution setting, as shown in Fig. 11. We also report results on the CLIC2020 test set at 768 resolution under DiffEIC’s protocol, i.e., resize & center-crop, 256 FID, as illustrated in Fig. 12. Across datasets, metrics, and evaluation settings, the OneDC consistently achieves SOTA performance, demonstrating strong robustness and generalization. The raw data used for evaluation is available at Table. 5 $\sim$  10.
We also report traditional pixel-level distortion metrics (PSNR and MS-SSIM [60]) to provide a more comprehensive analysis, as shown in Fig. 13. At extremely low bitrates, optimizing for PSNR often suppresses high-frequency details, resulting in blurred reconstructions [41]. While OneDC shows slightly lower PSNR than MS-ILLM, both perceptual metrics and qualitative examples clearly demonstrate its superior visual quality. Despite prioritizing perceptual quality, OneDC still achieves competitive PSNR compared to other diffusion-based methods. Moreover, on the MS-SSIM metric, OneDC matches MS-ILLM and outperforms all other baselines. These results confirm that OneDC delivers strong pixel-level fidelity alongside high perceptual realism, highlighting the overall effectiveness of our method.
Additional visual examples. We present more qualitative comparisons across four datasets: Kodak (Fig.16), CLIC2020—both full resolution (Figs.17) and $768\times 768$ cropped (Fig.18), and MS-COCO 30K (Fig.19). OneDC consistently outperforms prior SOTA methods, delivering superior visual quality across diverse content and resolutions, yet with the lowest bitrate cost.
Effectiveness of semantic distillation. To further evaluate the proposed semantic distillation strategy, we fine-tune the pretrained text-to-image one-step diffusion model [66], replacing its textual semantic condition with the hyperprior features produced by our semantic hyperprior decoder $h_{sem}$. This enables reconstruction using only the hyperprior signal. Fig. 20 presents additional reconstruction results on the COCO2017 validation dataset [35], extending Fig. 2 from the main paper. These results confirm that the distilled model captures richer semantic information, facilitating final reconstruction.
Bitrate allocation between hyperprior $\hat{z}$ and latent $\hat{z}$. We ablate the roles of hyperprior $\hat{z}$ and latent $\hat{y}$ on CLIC2020 test set by fixing the hyperprior budget to 0.0035 bpp and gradually increasing the bitrate of $\hat{y}$. Table 4 reports the results and reveals a clear division of each part’s role: (a) with zero bits allocated for latent $\hat{y}$, the codec still produces coherent reconstructions (also show in Fig. 20), indicating that $\hat{z}$ supplies a strong semantic information; (b) as bits are assigned to $\hat{y}$, fidelity and perceptual quality improve monotonically (e.g., LPIPS drops from 0.290 to 0.119), confirming that $\hat{y}$ is essential for encoding fine-grained details.
Table 3: Complexity analysis with model size. BD-Rate is calculated on the MS-COCO 30K dataset.
Model Params Enc. Time (s) Dec. Time (s) BD-Rate (Metircs)↓
LPIPS DISTS FID
MS-ILLM 181M 0.14 0.17 138.3% 253.0% 478.4%
DiffEIC 1.4B 0.32 12.4 305.0% 239.1% 341.0%
PerCo (SD) 3.8B+340M+955M* 0.58 8.80 538.8% 345.8% 59.6%
[…]
DDCM 950M - - - - -
OneDC 1.4B 0.15 0.34 0.00% 0.00% 0.00%
* Open-sourced PerCo includes an additional 3.8B BLIP2 caption model and 340M CLIP model.
Table 4: Bitrate allocation analysis on the CLIC2020 dataset.
Bpp $\hat{z}$ (ratio) Bpp $\hat{y}$ (ratio) Bpp Total PSNR↑ MS-SSIM↑ LPIPS↓ DISTS↓ FID↓
0.0035 (100%) 0.0 (0%) 0.0035 19.31 0.629 0.290 0.169 14.885
0.0035 (43%) 0.0047 (57%) 0.0082 23.13 0.790 0.163 0.089 6.223
0.0035 (27%) 0.0094 (73%) 0.0129 24.20 0.826 0.139 0.077 5.560
0.0035 (18%) 0.0157 (82%) 0.0192 25.25 0.856 0.119 0.068 4.979
Model size. We provide parameter counts, runtime comparisons (on 1024×1024 images), and BD-Rate results on the MS-COCO 30K dataset for better efficiency evaluation, as shown in Table 3. Compared to MS-ILLM, diffusion-based methods typically use larger models but achieve superior perceptual quality (e.g., lower BD-Rate with FID) due to stronger generative capacity. Unlike other diffusion-based codecs, OneDC avoids external caption models and multi-step sampling, enabling over 20× faster decoding while also achieving better rate-distortion performance.
Memroy usage. We also report memory usage under the resize & center-crop setting on $512\times 512$ Kodak images. PerCo (SD) requires about 22,220 MB of GPU memory, whereas DDCM uses 4,186 MB and OneDC uses 8,038 MB. These results highlight that incorporating a large language model, as in PerCo, substantially increases computational burden. DDCM achieves lower memory consumption by employing the SD model in a zero-shot manner. Although OneDC requires more memory than DDCM, its single-step design reduces inference cost compared to both multi-step PerCo and DDCM.
Figure 10: Comparison with original PerCo [9] on the MS-COCO 30K dataset.
Figure 11: Rate–distortion curves on the DIV2K dataset at full-resolution setting.
Figure 12: Rate–distortion curves on the CLIC2020 at resize & center-crop, 256 FID setting, with resized resolution 768.
Figure 13: Comparison of methods measured by PSNR and MS-SSIM at full-resolution setting.
Appendix B Training Details
Stage I training. This stage focuses on training the compression module and fine-tuning the one-step diffusion model [66] for the image reconstruction task. The training loss is defined as:
[…]
We use the L1 as the pixel-level loss and the LPIPS [24] as the perceptual-level loss. To support various bitrates, the rate-distortion trade-off parameter $\lambda$ is set to $\{0.6,1.0,1.8,2.9,4.6,7.4,12.2\}$. An auxiliary code prediction loss $L_{aux}$ is included with a weighting factor of $\alpha=0.001$. We train our model on the dataset introduced in [17]. Training is performed on 4×A100 GPUs for 800,000 steps, using a three-stage learning rate schedule with AdamW [38]: a) 5e-5 for the first 500,000 steps; b) 1e-5 for the next 200,000 steps; c) 1e-6 for the final 100,000 steps. During training, image patches of size $\{512,1024\}$ are randomly cropped with probabilities of $\{0.6,0.4\}$, respectively. The batch size is set to 32 for $512\times 512$ crops and 8 for $1024\times 1024$ crops (across 4 GPUs). This stage takes approximately 6 days as we use high-resolution patches for training.
Stage II training. This stage fine-tunes the one-step generator to improve reconstruction realism by better aligning the distribution of its outputs with that of real images. The training objective is:
$L_{stageII}=L_{distill}+\beta L_{recon}+\gamma L_{adv},\;\;\;\;\text{where:}$ (7)
$L_{distill}=\mathbb{E}_{t,\tilde{y}^{\prime}_{t}}[\epsilon_{fake}(\tilde{y}^{\prime}_{t},t)-\epsilon_{real}(\tilde{y}^{\prime}_{t},t)],\;\;L_{adv}=\mathbb{E}_{t,\tilde{y}^{\prime}_{t}}[-Disc(\epsilon_{fake}(\tilde{y}^{\prime}_{t},t),t)]$ (8)
Here, $L_{distill}$ represents the diffusion distillation loss [67], and $L_{adv}$ is the adversarial loss in the latent space, following [66]. $Disc$ denotes the discriminator network, which takes the mid feature in the diffusion U-Net as the input [66]. The variable $\tilde{y}_{0}$ is the latent output generated by the one-step diffusion model, and $\tilde{y}^{\prime}_{t}$ is its noised version at timestep $t$. We uniformly sample $t\in[20,640]$, since synthesizing high-frequency details does not require large noise levels. The weighting parameters are set as follows: $\beta=0.625$ balance the reconstruction and distillation terms, and $\gamma=0.001$ for the adversarial loss ($\gamma$ follows [66]).
Specifically, diffusion distillation [67] minimizes the expected Kullback-Leibler (KL) divergence between the time-dependent distributions of the target $p_{real,t}$ and the generator output $p_{real,t}$, thereby effectively transferring knowledge from the multi-step diffusion model to the one-step generator. The gradient used to update the one-step generator parameters $\theta$ is given by the difference between the score functions of the real and fake distributions:
$\nabla_{\theta}L_{distill}$ $=\mathbb{E}_{t,\tilde{y}_{0}}(\nabla_{\theta}\text{KL}(p_{fake,t}||p_{real,t}))$ (9)
$=-\mathbb{E}_{t,\tilde{y}^{\prime}_{t}}[(s_{real}(\tilde{y}^{\prime}_{t},t)-s_{fake}(\tilde{y}^{\prime}_{t},t))\frac{d\epsilon_{\theta}}{d\theta}]$ (10)
$=\mathbb{E}_{t,\tilde{y}^{\prime}_{t}}[(\epsilon_{fake}(\tilde{y}^{\prime}_{t},t)-\epsilon_{real}(\tilde{y}^{\prime}_{t},t))\frac{d\epsilon_{\theta}}{d\theta}]$ (11)
Here, $s_{\text{real}}$ and $s_{\text{fake}}$ are the score functions learned by multi-step diffusion model $\epsilon_{\text{real}}$ and $\epsilon_{\text{fake}}$ respectively. To ensure the fake score network $\epsilon_{\text{fake}}$ accurately tracks the evolving distribution of the one-step diffusion model, we update it using a standard denoising loss:
[…]
The improved version of diffusion distillation proposed in [66] introduces adversarial training in the latent space to further enhance distribution alignment. A discriminator is trained to differentiate between features extracted from real and generated images, using the following objective:
$L_{Disc}=\mathbb{E}_{t,x}[Disc(\epsilon_{fake}(\tilde{y}^{\prime}_{t},t))-Disc(\epsilon_{fake}(E_{\text{VAE}}(x)+n_{t},t))]$ (13)
Here, $E_{\text{VAE}}$ denotes the encoder of the diffusion model, and $n_{t}$ is the sampled noise in timestep $t$. The fake branch is then optimized with an additional adversarial loss:
$L=L_{fake}+\sigma L_{Disc}$ (14)
We follow [66] and set $\sigma=0.01$. During training, the fake branch and discriminator are each updated 10 times for every update of the one-step generator, ensuring stable adversarial optimization.
Training is conducted on 4×A100 GPUs for 1,000,000 steps. This stage requires around 12 days, as diffusion distillation introduces additional cost in addition to high-resolution training. The learning rate is fixed at 1e-6 (with AdamW) for the one-step generator, fake network, and discriminator. The batch size, cropping strategy and training data are identical to those used in Stage I.
Appendix C Model Details
The overall architecture is shown in Fig.14. To ensure better alignment with the latent space of the diffusion model, we extract features from its pretrained VAE encoder. Additionally, inspired by DiffEIC[33], we incorporate embeddings from the original input image to enrich the encoder with complementary spatial and semantic information. To jointly learn compressible latents and capture high-level semantics, we introduce a lightweight U-Net within the analysis transform. Its multi-scale design facilitates effective aggregation of both local textures and global structures. The extracted semantic features are integrated into the hyperprior branch, enhancing its representational capability.
For parameter-efficient adaptation, we insert LoRA [21] layers across all modules of the pretrained one-step diffusion U-Net, setting the LoRA rank to 64. The combined parameter count of the encoder and decoder components ($g_{a},h_{enc},h_{ctx},h_{sem}$ and $g_{s}$) is 394M, while the adapted diffusion U-Net contributes an additional 928M parameters (about 860M for SD1.5 and 68M for LoRA).
Further details of the semantic distillation mechanism are provided in Fig. 15. We adopt a Swin Transformer [37] to improve adaptability across different image resolutions during training.
Figure 14: Details of our model architecture. The pratrained VAE encoder is from SD 1.5. For the U-Net used in $g_{a}$, we use the implementation from the diffusers library [56].
Figure 15: Details of the auxiliary code predictor module. For Swin Transformer block [37], we set window size to 16, head dim to 64. W means normal window, SW means shifted window.
Appendix D Social Impact
Positive Aspects. Generative codecs offer substantial benefits by significantly reducing the storage and transmission demands of high-resolution imagery through the synthesis of perceptually convincing content. This improvement in efficiency can help broaden access to high-quality visual media, particularly in bandwidth-limited or resource-constrained environments. The resulting gains in data economy and speed of delivery have promising implications for applications in social communication and personal entertainment.
Negative Aspects. Despite these advantages, generative compression introduces synthesized content that may deviate from the original input, raising concerns about the fidelity and authenticity of reconstructed images. Ongoing research into multi-realism codecs [3] offers a potential path forward, enabling more transparent control over the trade-off between realism and fidelity.
Figure 16: Visual results on the Kodak dataset. The VAE-based MS-ILLM exhibits noticeable artifacts. Compared with our OneDC, the previous SOTA multi-step diffusion codec DiffC requires at least $1.4\times$ bitrate on these two examples. Zoom in for better view.
Figure 17: Visual results on the CLIC2020 test set. Zoom in for better view.
Figure 18: Visual results on the CLIC2020 test set ($768\times 768$). Compared with OneDC, the previous SOTA multi-step diffusion method DiffC requires $\geq 1.5$ higher bitrate while still producing slightly lower reconstruction quality on these examples. Zoom in for better view.
Figure 19: Visual results on the MS-COCO 30K dataset. PerCo, the previous best-performing method in terms of FID, requires $\geq 1.4$ bitrate over OneDC, but still results in suboptimal fidelity. Zoom in for better view.
Figure 20: Reconstructions from different semantic guidance. (a) Text prompts struggle to capture complex visual semantics, resulting in severe distortions when using a pretrained text-to-image one-step diffusion model [66]. (b) We finetune the model [66] for hyperprior-to-image generation. Hyperprior guidance yields more faithful reconstructions. (c) Our proposed semantic distillation further improves object-level accuracy.
Table 5: Compression quantitative evaluations, for the Kodak dataset with full-resolution.
Method BPP PSNR MS-SSIM LPIPS DISTS
MS-ILLM 0.0066 20.478 0.611 0.438 0.280
0.0156 22.224 0.722 0.292 0.177
0.0250 23.155 0.782 0.228 0.147
0.0447 24.731 0.846 0.158 0.135
0.0809 25.922 0.890 0.110 0.109
0.1535 27.532 0.928 0.073 0.081
0.2962 29.634 0.960 0.045 0.061
PerCo (SD) 0.0031 15.662 0.418 0.529 0.234
0.0324 19.344 0.673 0.299 0.162
0.1261 22.967 0.847 0.141 0.080
DiffEIC 0.0200 19.308 0.653 0.326 0.169
0.0375 20.970 0.741 0.242 0.134
0.0610 22.631 0.809 0.173 0.098
0.0916 24.147 0.861 0.129 0.077
GLC 0.0247 21.320 0.749 0.196 0.113
0.0286 21.729 0.768 0.180 0.104
0.0331 22.071 0.782 0.168 0.098
0.0374 22.279 0.791 0.161 0.095
DiffC 0.0063 18.494 0.545 0.450 0.201
0.0102 19.926 0.633 0.363 0.165
0.0155 21.212 0.705 0.288 0.139
0.0242 22.453 0.770 0.225 0.112
0.0423 23.823 0.832 0.163 0.087
0.0522 24.304 0.851 0.144 0.080
0.0625 24.704 0.865 0.130 0.074
0.0795 25.207 0.883 0.113 0.065
0.1227 25.949 0.905 0.091 0.055
OneDC 0.0034 17.141 0.472 0.380 0.204
0.0101 20.631 0.685 0.220 0.131
0.0165 21.513 0.741 0.183 0.110
0.0245 22.220 0.779 0.154 0.097
0.0354 22.979 0.817 0.133 0.085
0.0506 23.620 0.845 0.114 0.076
0.0775 24.396 0.877 0.096 0.066
0.1115 25.248 0.899 0.083 0.057
Table 6: Compression quantitative evaluations, for the CLIC2020 test set with full-resolution.
Method BPP PSNR MS-SSIM LPIPS DISTS FID
MS-ILLM 0.0045 22.815 0.731 0.365 0.216 44.954
0.0120 25.100 0.813 0.225 0.131 15.329
0.0195 26.275 0.853 0.173 0.106 8.748
0.0359 27.875 0.898 0.114 0.086 6.261
0.0654 29.229 0.928 0.080 0.066 4.481
0.1240 30.865 0.953 0.054 0.048 2.645
0.2377 32.833 0.972 0.035 0.035 1.660
PerCo (SD) 0.0022 15.340 0.473 0.589 0.326 76.658
0.0315 16.733 0.537 0.389 0.194 18.039
0.1249 17.419 0.554 0.306 0.091 5.013
DiffEIC 0.0142 20.152 0.739 0.308 0.185 18.427
0.0275 22.173 0.808 0.222 0.132 11.039
0.0487 24.656 0.865 0.153 0.089 7.278
0.0776 26.673 0.905 0.109 0.060 4.942
DiffC 0.0047 21.571 0.717 0.311 0.179 20.076
0.0079 22.914 0.771 0.253 0.144 12.332
0.0123 24.040 0.811 0.210 0.117 8.957
0.0197 25.241 0.849 0.171 0.092 6.813
0.0362 26.745 0.889 0.128 0.067 5.252
0.0556 27.772 0.911 0.103 0.053 4.414
0.0724 28.357 0.923 0.090 0.046 3.929
0.0872 28.733 0.930 0.082 0.042 3.685
OneDC 0.0035 19.305 0.629 0.290 0.169 14.885
0.0083 23.129 0.789 0.164 0.089 6.223
0.0130 24.201 0.826 0.139 0.077 5.560
0.0192 25.248 0.856 0.119 0.068 4.979
0.0279 26.096 0.879 0.103 0.060 4.234
0.0401 26.912 0.899 0.090 0.053 3.581
0.0617 27.804 0.921 0.076 0.044 2.977
0.0902 28.596 0.936 0.065 0.037 2.410
Table 7: Compression quantitative evaluations, for the MS-COCO 30K dataset at $512\times 512$ resolution.
Method BPP PSNR MS-SSIM LPIPS DISTS FID
MS-ILLM 0.0092 20.411 0.641 0.397 0.255 72.693
0.0196 22.708 0.766 0.257 0.169 17.992
0.0296 23.948 0.821 0.200 0.145 9.041
0.0488 25.517 0.875 0.138 0.122 4.100
0.0835 26.972 0.912 0.095 0.099 2.032
0.1510 28.734 0.944 0.063 0.075 0.990
0.2850 30.930 0.969 0.039 0.058 0.457
PerCo (SD) 0.0036 14.126 0.388 0.545 0.245 4.467
0.0329 18.124 0.676 0.311 0.159 2.748
0.1267 22.802 0.869 0.134 0.080 1.152
DiffEIC 0.0217 18.778 0.676 0.318 0.171 6.151
0.0407 20.564 0.770 0.229 0.133 3.929
0.0653 22.710 0.840 0.159 0.099 2.578
0.0975 24.606 0.886 0.116 0.077 1.911
DiffC 0.0083 16.170 0.495 0.524 0.271 90.989
0.0133 18.306 0.611 0.427 0.229 57.542
0.0198 20.159 0.705 0.334 0.187 28.909
0.0303 22.029 0.786 0.241 0.140 9.763
0.0514 24.170 0.859 0.155 0.095 2.805
0.0628 24.915 0.879 0.132 0.082 2.135
0.0745 25.520 0.894 0.115 0.074 1.809
0.094 26.264 0.910 0.098 0.064 1.518
OneDC 0.0034 16.083 0.460 0.376 0.211 13.607
0.0112 20.696 0.721 0.201 0.128 3.496
0.0179 21.924 0.779 0.166 0.111 2.817
0.0260 22.924 0.820 0.139 0.097 2.379
0.0371 23.809 0.852 0.118 0.086 2.044
0.0521 24.650 0.878 0.101 0.076 1.719
0.0789 25.674 0.905 0.082 0.064 1.365
0.1137 26.662 0.924 0.069 0.055 1.043
Table 8: Compression quantitative evaluations, for the DIV2K test set at full-resolution.
Method BPP PSNR MS-SSIM LPIPS DISTS FID
MS-ILLM 0.0054 20.647 0.647 0.396 0.234 79.455
0.0159 22.741 0.767 0.254 0.143 33.303
0.0262 23.837 0.822 0.198 0.116 22.428
0.0467 25.493 0.880 0.138 0.102 18.537
0.0837 26.878 0.917 0.097 0.080 13.574
0.1547 28.503 0.946 0.065 0.055 8.070
0.2895 30.545 0.969 0.040 0.040 5.123
PerCo (SD) 0.0022 14.529 0.381 0.570 0.320 95.345
0.0316 15.411 0.432 0.417 0.218 36.671
0.1249 15.340 0.405 0.355 0.105 11.314
DiffEIC 0.0182 18.615 0.663 0.327 0.190 34.527
0.0349 20.532 0.759 0.241 0.132 21.656
0.0575 22.604 0.831 0.172 0.089 14.026
0.0876 24.415 0.883 0.125 0.061 9.644
DiffC 0.0056 19.091 0.623 0.353 0.198 45.767
0.0096 20.345 0.695 0.291 0.158 32.086
0.0149 21.422 0.750 0.244 0.127 24.270
0.0238 22.599 0.803 0.199 0.098 18.541
0.0427 24.113 0.859 0.149 0.071 14.098
0.0640 25.158 0.890 0.120 0.056 11.859
0.0820 25.755 0.905 0.104 0.048 10.622
0.0976 26.143 0.914 0.096 0.044 9.938
0.1273 26.645 0.925 0.085 0.039 8.983
OneDC 0.0097 20.615 0.715 0.200 0.103 15.874
0.0161 21.719 0.771 0.168 0.086 13.680
0.0239 22.703 0.813 0.144 0.077 12.365
0.0345 23.546 0.846 0.124 0.067 11.242
0.0495 24.351 0.873 0.108 0.060 10.096
0.0749 25.257 0.901 0.090 0.050 8.325
0.1064 26.086 0.920 0.077 0.040 6.822
Table 9: Compression quantitative evaluations, for the Kodak dataset at resize & center-crop setting with $512\times 512$ resolution.
Method BPP PSNR LPIPS FID
MS-ILLM 0.085 25.683 0.11 36.08
0.159 27.296 0.072 28.556
0.304 29.395 0.044 24.448
PerCo (SD) 0.033 19.017 0.307 37.019
0.127 22.325 0.145 26.418
DDCM 0.03 22.066 0.222 32.031
0.038 22.551 0.19 29.117
0.05 23.013 0.161 25.647
0.095 23.606 0.138 24.215
0.149 24.069 0.124 23.199
OneDC 0.0109 20.259 0.223 36.862
0.0177 21.219 0.183 33.075
0.0257 21.970 0.154 30.947
0.0369 22.718 0.132 29.103
0.0524 23.363 0.113 27.428
0.0795 24.202 0.095 25.452
0.1139 25.021 0.082 23.448
Table 10: Compression quantitative evaluations, for the CLIC2020 test set at resize & center-crop setting with $768\times 768$ resolution.
Method BPP PSNR LPIPS FID
MS-ILLM 0.006 21.304 0.447 62.327
0.009 22.703 0.34 41.479
0.072 27.808 0.084 7.852
0.134 29.567 0.055 5.966
PerCo (SD) 0.003 15.339 0.517 30.409
0.032 19.018 0.269 12.869
0.126 23.387 0.122 5.419
DDCM 0.007 19.672 0.404 23.862
0.008 20.532 0.354 19.521
0.01 21.207 0.314 15.559
0.014 22.116 0.262 11.362
0.017 22.722 0.227 8.722
0.022 23.366 0.192 6.753
0.042 24.136 0.156 5.051
0.066 24.739 0.133 4.549
0.137 25.65 0.108 4.132
OneDC 0.0034 17.619 0.326 15.325
0.0098 22.378 0.172 7.953
0.0155 23.600 0.140 7.390
0.0228 24.641 0.117 7.011
0.0326 25.532 0.099 6.538
0.0461 26.416 0.084 5.874
0.0694 27.463 0.069 4.915
0.0993 28.471 0.058 4.046